### v3

# Graph RAG for Legal Norms: A Hierarchical, Temporal and Deterministic Approach

Abstract
This article proposes an adaptation of Graph Retrieval-Augmented Generation (Graph RAG) specifically designed for the analysis and comprehension of legal norms. Legal texts are characterized by a predefined hierarchical structure, an extensive network of references and a continuous evolution through multiple temporal versions. This temporal dynamism poses a significant challenge for standard AI systems, demanding a deterministic representation of the law at any given point in time. To address this, our approach grounds the knowledge graph construction in a formal, FRBRoo-inspired model that distinguishes abstract legal works from their concrete textual expressions. We introduce a multi-layered representation of Temporal Versions (capturing date-specific changes) and Language Versions (capturing linguistic variations). By modeling normative evolution as a precise sequence of these versioned entities, we enable the construction of a knowledge graph that serves as a verifiable "ground truth". This allows Large Language Models to generate responses based on accurate, context-aware, and point-in-time correct legal information, overcoming the risk of temporal inaccuracies. Through a detailed analysis of this formal Graph RAG approach and its application to legal norm datasets, this article aims to advance the field of Artificial Intelligence applied to Law, creating opportunities for more effective and reliable systems in legal research, legislative analysis, and decision support.
Keywords: Retrieval-Augmented Generation; Knowledge Graphs; Legal Tech; Legal AI; Legal Information Retrieval; Temporal Modeling; Knowledge Representation.
1 Introduction
In the contemporary era, where dependence on precise and accessible information is growing, the legal domain presents a unique challenge to professionals and analysis systems due to its vast volume and intrinsic complexity. Navigating this mass of data and efficiently extracting relevant information becomes a progressively arduous task. This difficulty not only burdens legal professionals but also exposes the limitations of conventional review and analysis methods – whether manual or digital – which are often slow and susceptible to errors. This complexity is aggravated by legal language itself, which, with its specific terminology and requirement for deep contextual understanding, presents essential nuances for rigorous interpretation and analysis.
The challenge is amplified by three defining characteristics of legal norms: the predefined hierarchical structure, the extensive cross-references (internal and external) and the multiplicity of temporal versions. In this scenario, Artificial Intelligence (AI) emerges as a promising field, offering tools and techniques capable of overcoming the barriers of traditional legal analysis and optimizing access to and comprehension of this vast knowledge.
Legal frameworks are inherently hierarchical. This structure manifests both between different types of norms (constitutions, statutes, regulations) and within individual legal texts themselves. Higher-level legal principles (e.g., constitutional provisions) frequently influence or restrict lower-level norms (e.g., specific regulations) – regarding the hierarchy of norms and legal principles, see [1]. On the other hand, legal norms are typically organized internally through grouping provisions such as titles, chapters, sections, and further subdivided into specific provisions like articles, paragraphs, items, clauses etc. Understanding the relationships across all levels of this hierarchy – both inter-norm and intra-norm – is essential for the interpretation and application of Law.
Additionally, the dynamic nature of legal norms, evolving through amendments, repeals, and new interpretations, results in multiple temporal versions. The ability to track and understand this evolution [2] is, therefore, essential for determining the applicable legislation at a given moment. This temporal dynamism presents a critical challenge for current AI systems, which, being inherently probabilistic, often cannot deterministically reconstruct the precise text of a law as it existed on a specific date. To ensure verifiable accuracy, our proposal is grounded in a formal, component-level versioning model, as detailed in [3], which provides a robust theoretical foundation for representing the diachronic evolution of legal texts.
These defining characteristics of legal norms, although representing significant analytical challenges, are indispensable for correct application and normative interpretation. Consequently, they demand analysis approaches that explicitly capture the structural and temporal dimensions of the legal norms, precisely the objective of the adapted Graph-based Retrieval-Augmented Generation (Graph RAG) methodology in this article.
Graph RAG emerges as an evolution of traditional RAG [4], which combines information retrieval with generation by Large Language Models (LLMs) [^1] LLMs (Large Language Models), such as GPT-4 or Gemini 2.0, are AI models trained to understand and generate natural language. For legal applications, see [5]. to enhance the relevance and precision of content. By enabling LLMs to base their answers on external knowledge sources, going beyond the data used in their training, RAG mitigates inherent limitations of these models, such as the propensity for hallucinations [^2] LLMs are trained on vast amounts of data but may lack specific or updated information. RAG addresses this problem by retrieving relevant documents and providing them as context to the LLM before it generates a response..
By utilizing knowledge graphs to represent and structure the relationships between concepts and entities extracted from these knowledge sources, Graph RAG [6] goes beyond traditional RAG by capturing the intricate and interconnected relationships within and between documents, which is particularly valuable for complex domains like Law. Legal knowledge is highly relational. Statutes refer to other statutes, clauses reference other clauses, and jurisprudence is based on precedents. Knowledge graphs are well-suited to represent these complex relationships.
Graph RAG presents significant potential for the analysis of legal norms. This technique can integrate structured information from knowledge graphs with unstructured textual content to create richer and more interconnected representations of legal knowledge. This integration allows capturing the complex relationships between legal entities, such as laws, doctrine, jurisprudence, concepts, and the temporal evolution of legal texts. By representing these legal entities, and their relationships, in a graph and linking them to the actual text of the legal norms and their different versions, modeled as a precise sequence of temporal expressions, Graph RAG can provide a more holistic understanding of Law [7].
The objective of this article is to contribute to the advancement of knowledge in Artificial Intelligence applied to Law, detailing the analysis of Graph RAG and its application to datasets of legal norms, with the goal of facilitating more effective systems to support legislative and legal work.
This paper is structured as follows. Section 2 first defines the core structural entities – Norm, Component, and Version – and introduces the formal model for their representation, which is essential for modeling legal documents, discussing their hierarchical relationships and the challenges of representing temporal changes. Section 3 then elaborates on our proposed adaptations to the Graph RAG framework, introducing the concept of Text Units, explaining the rationale for modeling temporal versions as aggregations rather than simple compositions, detailing the inclusion of essential metadata and alteration actions, and discussing strategies for structure-aware retrieval. Lastly, Section 4 concludes the paper by highlighting the key contributions and suggesting avenues for future investigation.
2 Structural Entities in the Graph
A Graph RAG system operates through the combination of several essential components. At its core lies a knowledge graph [8], which serves as a structured representation of knowledge. In this graph, nodes represent entities such as legal concepts, articles of law, or judicial cases, while edges symbolize the relationships between these entities, such as definitions, citations, or modifications. The knowledge graph provides a semantic foundation for understanding these connections. It explicitly models relationships, which are often implicit in unstructured text, making them accessible for retrieval and reasoning.
The initial step in both Graph RAG and traditional RAG approaches is to divide each text in the input dataset into segments, or “text chunks” - these segments are, throughout the process, converted into embeddings [9] and stored in a vector database [10]. Various segmentation strategies can be employed, such as sentence segmentation, paragraph segmentation, window-based splitting (according to the embedding model’s window size), and semantic segmentation (dividing text according to its semantic blocks), among others.
The second step in Graph RAG involves extracting Entities, along with their properties and relationships, from each text segment generated in the previous step. These extracted entities serve as the “building blocks” or nodes of the knowledge graph, with edges interconnecting them to represent the relationships derived from the normative texts.
In the context of legal norms, the entities extracted from the texts can be categorized into two groups:
  1. 1.
    Content Entities: traditional entities in the text of a legal norm such as types Person, Organization, Location, Event, Concept (metamodel entity [^3] Legal metamodel, in this context, refers to a conceptual model that defines the fundamental types of entities and relationships in the domain of Law, such as abstract legal concepts, principles, values, etc. Examples would be: ”Right to Life”, ”Principle of Legality”, ”Social Contract”.), Legal Norm (normative reference [^4] Normative reference represents a remission or citation to another legal norm, either in its entirety or to a specific provision.) etc.
  2. 2.
    Structural Entities
    • entities that represent the hierarchical structure of the text of a legal norm (its articulation), according to the formation rules of this type of text;
    • entities that represent the versions of this structure over time.
2.1 A Formal Model for Structural and Temporal Representation
To precisely model the structural and temporal complexity of legal norms, we ground our approach in the Functional Requirements for Bibliographic Records (FRBRoo) framework [11], adapted for the legal domain as detailed in [3]. This model distinguishes between abstract concepts and their concrete textual realizations, enabling a deterministic representation of legal text evolution. We define four primary types of structural entities:
  1. 1.
    Norm (as a Whole): Represents a legal norm as an abstract intellectual creation (e.g., the "1988 Brazilian Federal Constitution"). This corresponds to the Work (F1 Work) concept in FRBRoo, capturing the legal instrument’s identity independent of any specific text or amendment.
  2. 2.
    Component: Represents a hierarchical element within a norm (e.g., a title, chapter, or article) as an abstract concept. Each component is a Component Work, an identifiable part of the main Work that persists conceptually even as its text changes over time.
  3. 3.
    Temporal Version (TV / CTV): Represents the precise text of a Norm or Component as it exists at a specific point in time. A Component Temporal Version (CTV) is the TV of a specific Component. Each is a distinct, date-stamped entity, corresponding to the FRBRoo Expression (F2 Expression) concept.
  4. 4.
    Language Version (LV / CLV): Represents the text of a specific Temporal Version in a particular language. A Component Language Version (CLV) is a LV of a specific CTV. Each is also an Expression that derives from a single TV.
For the sake of clarity and to avoid terminological clutter, we will adopt a simplified notation for the remainder of this paper. The concepts that apply to a Norm as a whole (e.g., Temporal Version) and those that apply to its parts (e.g., Component Temporal Version) are structurally identical. Therefore:
  • We will use the term Temporal Version to refer to both the temporal version of a Norm and a Component.
  • We will use the term Language Version to refer to both the language version of a Norm and a Component.
The context will make it clear whether the entity refers to the norm as a whole or to one of its specific components.
The Norm and Component types exist in the conceptual Hierarchy dimension. The Temporal Version introduces the Time dimension [12], while the Language Version adds a Linguistic dimension.
For the sake of simplicity in the subsequent examples and diagrams, which focus on the Brazilian monolingual official context, we will often refer to a single Version entity. This entity should be understood as the primary Temporal Version and its corresponding Portuguese Language Version combined. This pragmatic simplification allows us to focus on the hierarchical and temporal modeling, while acknowledging that the underlying framework fully supports multilingual representations.
2.2 Legal Norm Structure in Practice
To structure the text of a Brazilian legal norm [13], the following components are defined [^5] In a more precise representation of the structuring of a legal norm, there would be the element Componente Autônomo Articulado (Autonomous Articulated Component), which would group all parts of the norm, including annexes. But, for didactic simplification, it will not be presented.. Each of these structural elements, such as Artigo or Capítulo, is modeled as a Component entity in our graph.
  • da Parte Inicial (of the Initial Part):
    • $\circ$
      Fórmula de Promulgação (Promulgation Formula): indicates the authority or institution responsible for creating the norm.
    • $\circ$
      Epígrafe (Epigraph): presents the full name of the norm.
    • $\circ$
      Ementa (Abstract/Summary): provides a concise summary of the norm’s content.
    • $\circ$
      Preâmbulo (Preamble): offers an introductory text that indicates the proposing body of the norm and, when applicable, its legal basis.
  • Articulação (Articulation):
    • $\circ$
      Encompasses the legal provisions, structured into Dispositivos [^6] A Dispositivo (Provision) of a legal norm is each element/component of the text that contains the prescription, i.e., the command, the rule of conduct, the permission, the prohibition, or the authorization that the norm establishes. It forms the normative core of the legal text. (Provisions) as illustrated in figures 1 and 2.
  • da Parte Final (of the Final Part):
    • $\circ$
      Fecho (Closing): details the place and date of issuance of the norm.
    • $\circ$
      Assinaturas e Assinaturas Complementares (Signatures and Complementary Signatures): include the necessary signatory information.
This hierarchical organization not only facilitates the systematic interpretation of legal texts but also enables precise navigation and retrieval of specific legal provisions.
The basic unit of a Brazilian legal norm’s articulation is the Artigo (Article). Each Article can be further subdivided into additional provisions (Components), as demonstrated in Figure 1.
Figure 1: Diagram of the hierarchical structure of possible subdivisions for the Article (Artigo) provision in Brazilian legislation. It shows Artigo (Article) $\to$ Caput (main section); Artigo (Article) $\to$ Parágrafo (Paragraph); Caput, Parágrafo (Paragraph) $\to$ Inciso (Item); Inciso (Item) $\to$ Alínea (Subitem); Alínea (Subitem) $\to$ Item (Sub-subitem).
Articles can also be organized into Dispositivos de Agrupamento (Grouping Provisions/Components), as illustrated in Figure 2.
Figure 2: Diagram of the hierarchical structure of grouping for the "Article" (Artigo) provision in Brazilian legislation. It shows Articulação (Articulation) $\to$ Parte (Part) $\to$ Livro (Book) $\to$ Título (Title) $\to$ Capítulo (Chapter) $\to$ Seção (Section) $\to$ Subseção (Subsection) $\to$ Artigo (Article).
As an example, Figure 3 presents the provision (Component) “Art. 12” of the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)), displaying its grouping provisions/Components (“Título II” and “Capítulo III”) and some of its child provisions (Components). Highlighted is the article, the basic unit of articulation. Annotations in orange are representing the type of each provision (Component).
Figure 3: Example of articulated text for Art. 12 of the Federal Constitution of Brazil (1988) with annotations indicating the types of hierarchical provisions/components.
2.3 Legal Norm Text Segmentation
In the original Graph RAG proposal, the first step involves segmenting the entire document into smaller text segments, or “text chunks”. In the subsequent step, each segment is processed—typically using a large language model (LLM)—to extract the identified entities, along with their properties and relationships, which are then added to the Knowledge Graph. However, when dealing with legal texts, the hierarchical structure is a fundamental characteristic that must be preserved to ensure accurate interpretation. Therefore, it is essential to extract and represent this hierarchical structure within the Knowledge Graph by mapping it to the previously defined structural entities: Norm (as a Work), Component (as a Component Work), and their corresponding Versions (as Expressions).
Ideally, semantic segmentation [^7] Semantic segmentation is a division of the document based on the meaning and structure of the text, rather than merely on superficial textual criteria such as sentences or paragraphs. should be employed to divide the document into text segments that correspond directly to its structural elements, as illustrated in Figure 3. In this approach, the initial segmentation step would simultaneously extract the structural entities, thereby capturing the hierarchical organization from the outset - the subsequent step would focus on extracting content entities, as outlined in the original Graph RAG proposal.
This proposed new first step (semantic segmentation oriented by the hierarchical structure of the legal norm text) could be done using an LLM (with fine-tuning or prompt engineering) prepared to recognize the structure of legal norms in the worked domain, generating a structured output in XML or JSON, for example. Another option would be to use a specialized Parser [^8] In the Brazilian case, an implementation of a Parser is available in the LexML project, which receives an unstructured text of a Brazilian legal norm and produces an XML file with the structured text. This parser, developed specifically for Brazilian legislation, represents a valuable tool for applying the proposed structured segmentation, as it automates the recognition and extraction of the hierarchical structure of norms, facilitating the creation of the Knowledge Graph. LeXML Brazil, Legislative Integration Project. Available at: https://projeto.lexml.gov.br. Accessed on: 15 Mar 2025. for segments of legal norm texts that recognizes and processes this structure – in both options, human review is recommended, given the complexity and importance of precision in interpreting the legal structure, and to ensure the correctness and quality of segmentation and extraction.
As an example, a passage of the the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)) received the semantic segmentation (like showed in Figure 3), producing the abstract entity of type Norm, represented by a circle in green, and those of type Component, represented in blue in the hierarchy presented in Figure 4.
For a more precise representation in the Graph, specialized entity types, derived from Norm, can be created for each type of norm in the treated legal system. In Brazil, there are, for example: "Constituição" (Constitution), "Emenda Constitucional" (Constitutional Amendment), “Lei Ordinária" (Ordinary Law), "Lei Complementar" (Complementary Law) etc. Thus, the entity representing a legal Work could be included in the graph using the specialized type.
Specialized entity types, derived from Component, can also be created for each type of provision found in the treated legal system. For example, in the Brazilian case, there are types like "Artigo" (Article), "Inciso” (Item), “Alínea" (Subitem), “Título” (Title), “Capítulo” (Chapter), presented in figures 1 and 2. Thus, the entity representing a Component Work could be included in the graph using the specialized type.
Figure 4: Illustration of hierarchical semantic segmentation, and typification of the structural entities, applied to a passage of the Federal Constitution of Brazil (1988), representing the Norm (Work) in green, Components (Component Works) in blue, and their associated "texts chunks".
2.4 Text Versioning over Time
As defined in our formal model, the textual evolution of a norm or component is captured through Temporal Version (TV / CTV) entities. Each TV / CTV represents the precise state of a text at a specific point in time and can be extracted from its original enactment or from a “compilação multivigente999“Compilação” (Compilation) in Brazil is a process in which, starting from the original version of a legal norm, all modifications commanded by subsequent norms are applied to its text, up to the specified date, thus producing an unofficial updated text of this norm on this date. It is ”Multivigente” (Multi-temporal) when all dates on which the norm underwent alteration are considered in the Compilation, producing an unofficial updated text (version) for each identified date. (multi-temporal compilation). Each TV / CTV, in turn, has at least one Language Version (LV / CTV) containing the text in a specific language.
A Component always has an original Temporal Version (CTV), which corresponds to the text (and label) enacted on the signature date of the original norm. Subsequent legislative acts (e.g., amendments) that alter, repeal, suppress, or renumber a component do not change the abstract Component (the Component Work) itself, but rather create a new Temporal Version for it. This new CTV captures the textual state of the component after the modification.
Since the full text of a norm is constructed by hierarchically grouping the texts of its individual components, the creation of a new CTV for any single component automatically generates a new CTV for its parent components, all the way up to the Norm as a whole (for the norm as a whole, its is a TV). This ensures that the compiled (unofficial consolidated) text of the entire legal norm is always consistent at any given date.
For example, Figure 5 illustrates the Version entities for the provision (Component) “Art. 6º” (and its caput - main section) of ’Constituição Federal brasileira’ (Brazilian Federal Constitution)" [^10] The structuring of the ”Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988) into ”Norm”, ”Components”, and ”Versions” can be visualized at https://normas.leg.br/?urn=urn:lex:br:federal:constituicao:1988-10-05;1988. Accessed on 15.03.2025.. In this monolingual example, each depicted "Version" entity represents both a Temporal Version and its corresponding Portuguese Language Version. The figure demonstrates how different versions, triggered by constitutional amendments, reflect changes over time.
Figure 5: Structure of the provision (Component) "Art. 6º" (and its caput - main section) of Brazilian Federal Constitution, illustrating four Version entities over time (light gray circles). Each represents a distinct Temporal Version (and its corresponding Portuguese Language Version) of the component, capturing the different wordings on different dates due to constitutional amendments.
2.5 Building the Graph
After semantically segmenting the text of a legal norm into its hierarchical structure, the elements of this hierarchy are inserted into the Knowledge Graph as nodes. As previously defined, we create nodes of type Norm (the Work) and Component (the Component Work), as illustrated in Figure 4. These nodes represent the conceptual backbone of the legal instrument, identified by properties like URN, name, and creation date.
The textual content itself is then linked to this abstract structure through Temporal Version and Language Version nodes. For each Component, we create:
  1. 1.
    A Temporal Version (CTV) node for each point in time its text was enacted or amended. This node is identified by a date-stamped URN.
  2. 2.
    A Language Version (CLV) node for each language in which that CTV’s text is available. This node is identified by a language-code-stamped URN.
Crucially, the text segment (text chunk) is associated with the Language Version node. This associated text is fundamental for the Retrieval-Augmented Generation process, as it represents the specific, contextualized content that will be retrieved and used by the LLM [14]. Figure 6 illustrates this multi-layered structure.
Figure 6: Representation of the multi-layered relationship in the graph: a Norm has Components, which have date-stamped Temporal Versions (CTVs), which in turn have language-specific Language Versions (CLVs). The Text Chunks are linked to the CLVs.
For example, the Norm (Work) node/entity “Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988), identified by urn: [^11] The formation of the urns presented in this work follows the standard proposed by the Brazilian LexML project. Available at: https://projeto.lexml.gov.br. Accessed on: 15 Mar 2025.
It has a Component (Component Work) node/entity “caput do Art. 6º” (caput is the main text of the article) identified by:
This component has a original Temporal Version (Expression)—the textual formulation of the component on the date of norm’s signature:1988-10-05—is identified by adding @YYYY-MM-DD:
And this TV has a primary Language Version (Expression) for Portuguese ("pt"), identified by:
This Language Version node is associated with the Text Chunk containing the original wording (in Portuguese) of the caput of Art. 6º, as illustrated in Figure 5:
"São direitos sociais a educação, a saúde, o trabalho, o lazer, a segurança, a previdência social, a proteção à maternidade e à infância, a assistência aos desamparados, na forma desta Constituição."
("Social rights include education, health, work, leisure, security, social security, protection of motherhood and childhood, and assistance to the destitute, as prescribed by this Constitution.")
Unlike static documents, the text of a legal component evolves. This evolution is represented by a set of Temporal Versions, each capturing a state of the text at a specific time. Figure 7 displays the four Temporal Versions of the caput of Art. 6º, with each CTV linked to its corresponding Portuguese CLV and text chunk—are displayed: one original CTV and three resulting from changes (text change) promoted by “Emendas Constitucionais” (Constitutional Amendments).
For a more precise representation in the Graph, specialized entity types can be derived from Temporal Version to represent the nature of the change found in the treated legal system. For example, in Brazil, there are types like "Original Version”, “Alteração de Texto” (Text Change Version), “Renumeração” (Renumbering Version), “Revogação” (Repeal Version) etc. Thus, the entity representing the temporal version could be included in the graph using the specialized type.
Figure 7: Representation of multiple Temporal Versions for the component (…!art6_cpt). Each CTV has a corresponding Language Version (in Portuguese) which is linked to its unique Text Chunk, illustrating the component’s textual evolution over time.
The core proposal of this work is to build a knowledge graph that models this complete hierarchical, temporal, and linguistic structure. It is important to note that text segments (text chunks) are linked to Language Version entities, not to Component or Temporal Version entities. This ensures that retrieval is always precise to a specific text in a specific language at a specific time.
This detailed representation allows answering questions about the text of a component with full temporal context: for example, if there was any change in the text over time; what changes it underwent; or what the text was on a specific date, and in which language.
3 Adapted approach
3.1 Text Units
Regarding the text segment (text chunk) associated with each Language Version (LV) entity, in cases of subdivisions of a caput, such as “parágrafos” (paragraphs) and enumeration elements (“incisos” (items) and “alíneas” (subitems)), we suggest adopting the strategy proposed by [15]. In this strategy, the text segment to be considered for a component is not only the text directly resulting from the semantic segmentation of the norm but the semantically complete text of the component, derived from post-segmentation (and post-compilation) processing. The cited author recognizes the hierarchical and interdependent nature of legal norm components, where, for example, the meaning of an enumeration component (e.g., “inciso” (item)) is frequently dependent on the context of its parent component (e.g., the caput).
In other words, when including in the graph the original Language Version of the provision "inciso I do caput do Art. 5º” (item I of caput of Art. 5), the text linked to this entity will not be just:
"I - homens e mulheres são iguais em direitos e obrigações, nos termos desta Constituição;"
("I - men and women are equal in rights and obligations, under the terms of this Constitution;")
but rather the text formed by the content of the article’s caput’s CLV (from the same Temporal Version) plus the content of the item itself. This can be visualized in Figure 8, where the parent’s text was placed as a prefix in square brackets.
"[Todos são iguais perante a lei, sem distinção de qualquer natureza, garantindo-se aos brasileiros e aos estrangeiros residentes no País a inviolabilidade do direito à vida, à liberdade, à igualdade, à segurança e à propriedade, nos termos seguintes:] I - homens e mulheres são iguais em direitos e obrigações, nos termos desta Constituição;”
("[Everyone is equal before the law, without distinction of any kind, guaranteeing to Brazilians and foreigners residing in the Country the inviolability of the right to life, liberty, equality, security, and property, under the following terms:] I – Men and women are equal in rights and obligations, under the terms of this Constitution;")
Figure 8: Example of contextual (cumulative) Text Units for the Language Versions of the "incisos’ (items) of caput of Art. 5, demonstrating the inclusion of the parental caput text to provide complete semantic context.
Conversely, the Language Version of an article’s caput will have as its linked text its own text accumulated with the text from the CLVs of its child components (all belonging to the same Temporal Version), as exemplified in Figure 9. The same logic applies recursively up the hierarchy to articles, groupers, and the norm as a whole.
According to the cited author, this would bring more “aboutness” [^12] The term ”aboutness” refers to the capacity of a text to faithfully represent its central theme. For theoretical discussions, see [17]. to the texts linked to the entities [16]. In other words, the cumulative “Text Unit”, by including the textual context of parent or child components, represents the central theme and legal meaning of the component in question more completely and comprehensively.
Figure 9: Example of a cumulative Text Unit for the Language Version of the caput of Art. 5 of the "Constituição Federal Brasileira de 1988" (Brazilian Federal Constitution of 1988).
Considering this new approach, we will call the texts associated with Language Version entities Text Units, instead of Text Chunks. They are not necessarily the pieces of text resulting directly from the initial segmentation of the norms’ texts, but rather the contextually complete content of each component, ready for embedding and retrieval.
3.2 Temporal Versions as Aggregations
When a subsequent norm, on date D1, commands an alteration to the content of a component C, a new Temporal Version (CTV) is created for that component. This is illustrated in Figure 10, which shows a component’s lifecycle as a series of date-stamped CTVs.
Figure 10: Temporal Versions of component C over time. Each node represents a distinct textual state of the component on a specific date.
Considering the proposal of Text Units resulting from the accumulation of text in the hierarchy, consequently, a new Temporal Version CTV (and its related CLV) must also be produced for every component of norm N that:
  1. 1.
    groups, directly or indirectly, component C, including the norm as a whole (as level 0; in this case, it is a TV);
  2. 2.
    has its Text Unit formed by the accumulation of the text of C as a prefix (for example: "parágrafos” (paragraphs), “incisos” (items) and “alíneas” (subitems)).
This propagation of change is where the concept of aggregation becomes critical. Consider Figure 11. On the norm’s original date (1988-10-05), the original CTV of the component "tit2" is a composition of the original CTVs of all its children from that same date.
On 1993-09-14, a new CTV for component “tit2 cap4” was created due to an amendment. A naive approach would be to create new 1993-09-14 CTVs for all other children of "tit2" (even the unchanged ones) and then compose them to form a new CTV for "tit2". This would be a Composition [^13] In UML terms, Composition represents a strong whole-part relationship where the ”part” belongs exclusively to one ”whole” and shares its lifecycle; if the whole is deleted, the part is too. In our initial comparison, this would mean a parent version strictly owns child versions from the exact same date.. This approach is inefficient, creating redundant data and obscuring which components actually changed.
Our proposal, however, is to model the new 1993-09-14 CTV for "tit2" as an Aggregation [^14] Aggregation, in UML, represents a weaker whole-part relationship where the ”part” can exist independently of the ”whole” and can potentially be shared by multiple wholes. In our proposed model, this reflects how a child version (the ”part”) can be reused by multiple parent versions (the ”wholes”) across different times without its existence depending on any single one. of its children’s most recent CTVs. This means the new CTV for "tit2" is formed by the new CTV of "tit2 cap4" and the reused, pre-existing CTVs of all its other, unaltered children.
This aggregation logic is shown in Figure 11. New CTVs for "tit2" are created only when one or more of its children are amended, and these new parent CTVs simply point to the latest available CTV of each child.
Figure 11: New Temporal Versions of the component "tit2" derived from new CTVs of some of its children. Unchanged child components have their most recent CTV reused (e.g., the 1988-10-05 CTV of tit2_cap2 is aggregated into the 1993-09-14 CTV of tit2).
This model establishes that a child’s Temporal Version does not "belong" exclusively to a single parent CTV but can be aggregated into multiple parent CTVs across different points in time, as illustrated in Figure 12. When retrieving the state of a component Cx for a date Dn for which no CTV exists, the system simply traverses back in time to find the most recent CTV prior to Dn. It is exemplified by the Version V1 of component C3, which contributes to the versions V1 and V2 of component C1, demonstrating that a child version can participate in multiple parent versions in different temporal contexts. The orange arrows indicate the aggregation relationship.
Figure 12: Diagram of aggregation relationships between Temporal Versions, illustrating how child CTVs can be reused by multiple parent CTVs at different times.
This representation of Temporal Versions as aggregations ensures uniqueness and efficiency. Instead of creating a complete version of the entire document tree for each amendment date, we only create new CTV nodes for components whose text (or whose descendants’ text) has actually changed. This eliminates redundancy, provides a non-ambiguous and economical representation, and accurately reflects the real-world legislative process.
In the context of RAG, this allows the system to retrieve not only the specific CTV (and its CLV/Text Unit) of a component, but also the precise context of its parent and child CTVs for any given moment, facilitating complex queries about the temporal evolution of norms [2].
3.3 Modeling Normative Events as Actions
The lifecycle of a Temporal Version (TV/CTV) — its creation and termination — is governed by normative events, such as amendments, repeals, or original enactments. To capture this causality explicitly, we introduce an Action entity into our graph. This approach is grounded in event-centric ontologies for the legal domain, which model legislative changes as discrete Events that occur at a specific time and trigger consequences [18].
An Action node represents a specific command within a legal norm that affects another norm or component. Examples include text amendments, repeals, suppressions, or renumberings. As illustrated in Figure 13, an Action has a dual role:
  1. 1.
    It terminates the validity of a prior Temporal Version.
  2. 2.
    It produces a new Temporal Version.
It can link the CTV (and the component) of the amending norm (the cause) to both the old TV/CTV that it terminates and the new TV/CTV that it creates.
Figure 13: Representation of a legislative Action (Event) in the knowledge graph. The Action, commanded by the Art.1’s caput of the Brazilian Constitutional Amendment 26, terminates the validity of the original 1988-10-5 CTV of Art. 6’s caput of the Brazilian Federal Constitution of 1988 and produces its new 2000-02-14 CTV.
For a more precise representation, specialized entity types derived from Action can be created, such as "Text Change Action" or "Repeal Action".
To make these events semantically retrievable, we propose creating a descriptive Text Unit for each Action. This Text Unit is a natural language summary of the action, its cause, and its effects. For the action illustrated in Figure 13, the content of this new Text Unit would be:
"A Emenda Constitucional nº 26, de 14 de fevereiro de 2000, por meio do caput do seu Art. 1º, deu uma nova redação ao caput do Art. 6º da Constituição Federal de 1988. Essa alteração encerrou em 2000-02-14 a vigência da versão original desse dispositivo (de 1988-10-05) e instituiu uma nova versão com vigência a partir de 2000-02-15, cujo texto passou a ser: ’São direitos sociais a educação, a saúde, o trabalho, a moradia, o lazer, a segurança, a previdência social, a proteção à maternidade e à infância, a assistência aos desamparados, na forma desta Constituição.”  [^15] The validity began on 2000-02-15 because the Amendment was published on that date, and its text stipulated: ’This Constitutional Amendment enters into force on the date of its publication.
("Constitutional Amendment no. 26, of February 14, 2000, through the caput of its Art. 1º, provided a new wording for the caput of Art. 6º of the Brazilian Federal Constitution of 1988. This alteration terminated on 2000-02-14 the validity of the original version of this provision (from 1988-10-05) and established a new version effective from 2000-02-15, whose text became: ’Social rights include education, health, work, housing, leisure, security, social security, protection of motherhood and childhood, and assistance to the destitute, in the manner prescribed by this Constitution.’")
Although this alteration command is implicitly present in the text of the amending norm (Figure 14), generating a dedicated Text Unit for the Action entity makes the event’s details explicit and directly queryable. It structures and completes the information about the command, clarifying, for example, the validity periods of the affected TV/CTVs.
Figure 14: Text of the Emenda Constitucional nº 26 de 14/02/2000 (Constitutional Amendment no. 26), which contains the command for the alteration action.
This generated Text Unit is then associated with the Action entity in the graph. Since this content is a linguistic description, it can exist in multiple languages, with each translation being a distinct Text Unit of the same Action node.
Like other Text Units, it receives its own embedding. This allows for sophisticated semantic searches not only on the content of legal norms (CLVs Text Units) but also on the events that shape their evolution (Actions Text Units), facilitating the traceability of normative alterations and improving the understanding of the life cycle of legal norms in the context of RAG [19].
Furthermore, these Action Text Units are also crucial for enriching the response generated by the LLM in the RAG process. When the system retrieves a specific Language Version as relevant context, the Text Unit of the Action that created or terminated its corresponding Temporal Version can also be provided to the LLM. This gives the model valuable information about the history and validity of the component, allowing for more complete and contextually rich responses. For instance, the LLM could then generate an answer that not only provides the text of a provision but also explains when it came into force, when it was repealed, and by which specific legislative act it was altered.
3.4 Metadata as Text Units
Besides the textual content of legal norms captured in Language Versions, other sources of information can be used to feed the knowledge graph. A proposed technique is to generate Text Units from structured metadata [^16] Metadata is data about data. In this context, structured or textual information describing the properties and characteristics of legal norms, such as publication date, authorship, type of norm, relationships with organizations, etc.. These Text Units are textual summaries associated directly with the abstract Norm and Component entities (the Works). This is similar to the entity summaries in the original Graph RAG proposal, but with a key difference: this information is generally not contained within the norm’s text itself and must be sourced externally or from graph properties.
For example, for the Norm entity representing the legal work “Lei nº 10.406 de 10/01/2002”, a metadata Text Unit would be generated to textually describe its properties [^17] This Text Unit was generated from the structured metadata registered in the knowledge graph of the ”Sistema de Gestão de Normas Jurídicas (Sigen)” (Legal Norm Management System) of the Federal Senate of Brazil. This metadata can also be viewed publicly, in structured form, through the Normas.leg.br portal, as in the example: https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406. Accessed on 2025-03-22. - like any textual description, this Text Unit can be generated in multiple languages.
"São apresentados a seguir os metadados gerais da norma jurídica federal ’Lei nº 10.406 de 10/01/2002’: é do tipo ’Lei Ordinária’; está no nível ’Legal’ do ordenamento jurídico brasileiro; possui como urn: ’urn:lex:br:federal:lei:2002-01-10;10406’; possui como URL principal (uri): ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; possui como URL alternativa: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’; foi apreciada e aprovada pelo ’Congresso Nacional’; foi assinada em Brasília, 10 de janeiro de 2002; 181º da Independência e 114º da República; foi assinada por FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; foi oficialmente publicada pela ’Imprensa Nacional’, tendo como data de publicação original e oficial: 11/01/2002; tem como nomes alternativos: ’LEI-10406-2002-01-10’ e ’Código Civil (2002) (CC)’; tem como ementa: ’Institui o Código Civil’."
("The general metadata for the federal legal norm Law no. 10,406 of 2002-01-10 are presented below: it is of the type ’Ordinary Law’; it is at the ’Legal’ level of the Brazilian legal system; its URN is: ’urn:lex:br:federal:lei:2002-01-10;10406’; its main URL (URI) is: ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; its alternative URL is: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’; it was reviewed and approved by the ’National Congress’; it was signed in Brasília, on January 10, 2002; 181st year of Independence and 114th year of the Republic; it was signed by FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; it was officially published by the ’National Press’, with the original and official publication date being: 2002-01-11; it has alternative names: ’LEI-10406-2002-01-10’ and ’Civil Code (2002) (CC)’; its summary is: ’Institutes the Civil Code’.")
Informative relationships between legal norms, such as regulation, succession, or correlation, are also modeled as relationships between Norm entities (Works). These relationships do not generate new temporal versions but are fundamental for semantic retrieval. We can therefore generate Text Units to describe them.
Considering the legal norm ’Constituição do Brasil (1967)’ (Brazilian Constitution of 1967), the following are the Text Units [^18] It is not necessary to generate a Text Unit for the other side of the relation, since an LLM and a RAG retrieval already semantically understand this reversal (”succeeded” vs “was succeeded by”, for example). for its succession relationships, as shown in Figure 15.
Figure 15: Example representation of inter-normative relationships (succession) between different Norm (Work) entities in the knowledge graph.
A norma jurídica ’CONSTITUIÇÃO DO BRASIL (1967)’ sucedeu a norma jurídica ’CONSTITUIÇÃO DOS ESTADOS UNIDOS DO BRASIL (1946)’
(The legal norm ’Constitution of Brazil (1967)’ succeeded the legal norm ’Constitution of the United States of Brazil (1946)’.)
A norma jurídica ’Constituição da República Federativa do Brasil’ sucedeu a norma jurídica ’CONSTITUIÇÃO DO BRASIL (1967)”
(The legal norm ’Constitution of the Federative Republic of Brazil’ succeeded the legal norm ’Constitution of Brazil (1967)’)
Text Units can also be created for relationships involving Component entities, for example:
"O dispositivo ’Art. 5º, caput, XII –’ da ’Constituição da República Federativa do Brasil’ é correlato da norma jurídica ’Lei nº 14.063 de 23/09/2020’."
("The provision ’Art. 5º, caput, XII -’ of the ’Constitution of the Federative Republic of Brazil’ is correlated with the legal norm ’Law no. 14,063 of 2020-09-23’.")
This approach aligns with the concept of “Multi-aspect embeddings[20], which proposes using multiple vectors to represent different aspects of an entity. A Norm or Component (Work) can have multiple associated Text Units (including, in various languages):
  • One Text Unit for its intrinsic properties (metadata summary).
  • A separate Text Unit for each of its relationships (e.g., succession, correlation).
Thus, instead of a single vector, an abstract entity like a Norm will be associated with multiple vectors/points in the vector space. Each vector captures a different "lens" on the entity, enabling more contextualized and granular semantic retrieval.
These metadata Text Units can be generated externally or from the structured properties and relationships already present in the graph [^19] As these properties (publication date, alternative names, level in the legal system, etc.) and relationships (correlation, succession, etc.) generally cannot be directly derived from the texts of the legal norms, they must be populated in the graph either manually or through semi-automatic extraction processes using other information sources.. Converting them to a textual format allows for their uniform processing within the Graph RAG workflow.
Each metadata Text Unit will have its own embedding. Consequently, vector search can retrieve information about a Norm or Component based on its metadata, even if that information is not present in its textual content. Figure 16 illustrates this concept, showing how Text Units can be derived both from Language Versions (content of the compiled legal norm) and from the abstract Norm/Component (and Action) entities themselves (metadata).
Figure 16: Knowledge graph illustrating how Text Units are derived from two sources: from Language Versions (representing content) and from the abstract Norm/Component (and Action) entities (representing metadata).
3.5 Hierarchical Grouping as Communities
The hierarchical structure of a legal norm provides a natural, intrinsic form of community organization. The grouping elements described earlier—such as Titles, Chapters, and Sections—are not merely textual containers; they are abstract Component entities that provide thematic context. As observed in Figure 4, the Component "Art. 5º" is part of the "CAPÍTULO I" Component, which in turn is part of the "TÍTULO II" Component. This creates a nested hierarchy of Works that ascends to the main Norm.
This structure directly aligns with the concept of "communities" in Graph RAG, where communities are densely interconnected clusters of nodes. In our model, the predefined legal hierarchy is a primary, curated form of community structure, superior to purely algorithmic community detection.
Beyond this internal hierarchy, legal norms are often classified into external, cross-cutting themes [^20] Brazilian federal norms are classified according to the “Sistema de Classificação de Normas Jurídicas do Congresso Nacional” (System for Classification of Legal Norms of the National Congress), with examples of themes including: “Children and Adolescents”, “Extraordinary Credit”, “Indigenous Population”, “Basic Sanitation”, ’General Social Security System’, etc.. These themes can be modeled as high-level community Theme entities (nodes) that group multiple Norm entities (Works) on cohesive subjects, creating inter-norm communities, as exemplified in Figure 17.
Figure 17: Diagram illustrating inter-norm aggregation by legal Theme entities/nodes (orange), represented as higher-level communities that group Norm (Work) entities (green) in the knowledge graph.
This thematic classification can also be applied at a more granular level, linking Theme nodes directly to specific Component entities, as shown in Figure 18. This allows for a richer, multi-faceted classification of legal provisions beyond their formal position in the document hierarchy.
Figure 18: Diagram illustrating associations between Component (Work) entities and legal Theme (Community).
These Theme nodes will be inserted into the Knowledge Graph, each with an associated Text Unit describing the theme’s scope (there can be a Text Unit for each language). This description could be produced by a human domain expert or generated semi-automatically with LLM assistance, where the LLM produces a summary based on the definitions of sub-themes or the content of the associated norms. The latter approach would require careful specialized human review, as the quality of the summary is critical. In either case, using predefined thematic structures provided by official sources is preferable, as it ensures the semantic grouping reflects established legal practice.
3.6 Structure-Aware Retrieval
A key advantage of our graph-based model is enabling structure-aware retrieval. By representing the full hierarchy of Themes, Norms, Components, and their Versions, the user can precisely define the scope of their query. For example, a user could choose to search within:
  • A specific Theme, retrieving Text Units from all sub themes, norms and components classified under it.
  • A specific Norm (e.g., the entire Civil Code), searching across all its components.
  • A specific Component (e.g., a single Title or Article), restricting the search to its subdivisions.
  • A specific Temporal Version of a norm or component, querying only the text as it existed on a particular date.
This process is illustrated in Figure 19.
Figure 19: Diagram illustrating how a user can select a scope (e.g., a Theme, Norm, Component or Version) to filter the retrieval of relevant Text Units from the knowledge graph.
Choosing a scope acts as a powerful filter on the set of Text Units to be considered for retrieval. Instead of performing a vector search across the entire legal corpus, the system first traverses the graph from the selected entry point (e.g., a "Title" Component) to identify your own Text Units and those of all your descendants. Only the Text Units belonging to this filtered subset are then used in the semantic search.
This hierarchical filtering allows for more precise "semantic navigation". The user can move between different levels of detail, from a broad thematic query to a focused analysis of a single article’s evolution. This structured, hierarchy-guided exploration of legal knowledge is more efficient and yields more relevant results than a flat search across an entire, undifferentiated database of text chunks.
4 Conclusion and Future Work
Aiming to enhance the analysis and comprehension of legal norms in the era of Artificial Intelligence, this article presented a detailed proposal for the construction of a legal knowledge graph. By grounding the graph’s design in a formal, FRBRoo-based model for temporal and linguistic versioning, we have outlined a robust methodology for representing the complex lifecycle of legal texts. The central objective was to expand the scope of queries regarding legal norms, allowing AI systems to address questions ranging from literal text retrieval to complex inquiries requiring a deterministic understanding of the hierarchical structure, intrinsic semantics, and temporal evolution of law.
In summary, the main contributions of this work include:
  • A formal, multi-layered model for legal knowledge graphs, distinguishing between abstract Works (Norms and Components) and their concrete textual Expressions (Temporal and Language Versions). This provides a solid theoretical foundation for representing legal documents.
  • The modeling of Temporal Versions as Aggregations, reflecting hierarchical interdependence and enabling an accurate and efficient representation of normative evolution by reusing unchanged components over time.
  • The proposal to generate specific Text Units for normative Events (Actions), making the norm-modifying events themselves and their effects (e.g., a change in validity) semantically retrievable.
  • The generation of Text Units from structured Metadata associated with abstract Norm and Component entities, allowing non-explicitly textual information to be included in vector search and aligning with multi-aspect embedding concepts.
  • The interpretation of the intrinsic legal hierarchy and extrinsic thematic classifications as curated Communities within the graph, enabling more targeted retrieval strategies.
  • The proposal of Structure-Aware Retrieval, which leverages the graph’s hierarchy to allow users to precisely define a search scope, refining the precision and relevance of responses.
  • A Graph RAG approach that seeks to minimize creation costs by prioritizing pre-existing, curated structures (both hierarchical and thematic) over computationally intensive, purely algorithmic methods.
The main innovation lies in the detailed adaptation and expansion of the Graph RAG paradigm for the legal field. By integrating a formal versioning model inspired by library science (FRBRoo), we provide a deterministic "ground truth" that is essential for reliable AI in a high-stakes domain like Law. This structured approach, combining a formal hierarchy, explicit versioning, and curated thematic communities, provides a rich and natural contextual basis for legal RAG strategies.
The proposal to use the hierarchy of communities and themes designed or reviewed by domain experts reflects the original intent of legislators and ensures informational curation, an essential aspect for the reliability of AI systems in Law, in addition to minimizing the need for complex and costly graph construction processes.
For future work, several research possibilities open up:
  • It is fundamental to empirically validate the effectiveness of the proposed Graph RAG approach through experiments with legal norm datasets and rigorous evaluations of the quality and precision of the responses generated by LLMs.
  • Although the proposal of this work significantly advances the representation of normative complexity, it primarily focused on modeling the hierarchical structure and temporality. The deep extraction and semantic integration of Content Entities (people, organizations, legal concepts etc.) and their complex interrelationships [7] remain a challenge reserved for future investigations.
  • Investigating the use of Artificial Intelligence techniques, such as language models and clustering algorithms, to support human curation in detecting thematic communities and suggesting hierarchical structures represents another promising direction.
  • Exploring the practical application of legal Graph RAG in legal decision support systems, legislative impact analysis tools, and platforms for facilitated access to legal information for legal professionals and citizens can demonstrate the concrete value and feasibility of the approach in real-world and diverse scenarios.
  • Integrating Bigraphs [21] with the proposed Graph RAG model could further enhance the representation of spatio-temporal relationships and enable dynamic memory operations, such as automatic version consolidation or the identification of normative conflicts via semantic link analysis.
  • Investigating the applicability and potential adaptation of the proposed Graph RAG approach beyond legal norms to other fundamental types of legal documents, such as jurisprudence (case law) and contracts. While the current focus leverages the predefined hierarchy of norms, exploring how this framework could be extended or modified to handle the different structural conventions (e.g., case structure, contract clauses) and temporal dynamics (e.g., precedential relationships, contract amendments) inherent in these other domains would be a valuable direction. This analysis would help clarify the approach’s scope of generality and its limitations, further contextualizing its contribution within the broader field of Legal AI.
Ultimately, by integrating structure, cross-references, temporality, and semantics into a legal knowledge graph, this work offers a step towards more intelligent, transparent, and effective Artificial Intelligence systems for the complex and fundamental domain of Law. By facilitating precise information retrieval and making legislative analysis more accessible through a navigable and contextualized representation, this approach has the potential to democratize access to legal information and empower professionals and citizens in an increasingly regulated world.